Initially a soccerball is going 37.5m/s south. In the end, it it travelling at 2.6m/s south. The balls changing in momentum is 17.24kg m/s North. Find the ball's mass.
1
Expert's answer
2020-04-20T10:42:21-0400

The balls changing in momentum is given by

"\Delta {\bf p}=m({\bf v_f}-{\bf v_i})"

In projection on the North direction

"\Delta p=m(v_i-v_f)"

Hence, the mass of ball

"m=\frac{\Delta p}{v_i-v_f}=\frac{17.24}{37.5-2.60}=0.5\:\rm kg."